OpenStudy (anonymous):

If 0 degrees ≤ A ≤ 180 degrees, and tan A=-√3, how do I find the value of sinA?

OpenStudy (anonymous):



OpenStudy (anonymous):

why is √3 the opposite and 1 the adjacent? ahh..

OpenStudy (anonymous):

there is an angle \(A\) with \(\tan(A)=-\sqrt{3}\) find the hypotenuse by pythagoras and then take the usual ratios

OpenStudy (anonymous):

to ask is to answer, yes?

OpenStudy (anonymous):

ohh, yes.

OpenStudy (anonymous):

hypotenuse is \[\sqrt{\sqrt{3}^2+1^2}=\sqrt{4}=2\] and all should be good from there. just keep in mind that you are in quadrant II so sine is positive but cosine is negative

OpenStudy (anonymous):

ohhh.. alright. By the way, do I have to memorize special triangles and their angles? For ex. sin60 degrees = √3/2.

OpenStudy (anonymous):

it certainly wouldn't hurt. they are all on the cheat sheet i am attaching. take a look

OpenStudy (anonymous):

ohh, thank you!

OpenStudy (anonymous):

yw